# The currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters

### v1

@@
22email: sami.dib@gmail.com
@@
We categorized clumps, embedded clusters, and open clusters, and conducted a comparative analysis of their physical properties. Overall, the radii of open clusters are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters. The open clusters have significantly larger masses than embedded clusters, by about one order of magnitude. Given the current mass distribution of clumps in the Milky Way, the evolutionary sequence from a single clump evolving into an embedded cluster and subsequently into an open cluster cannot account for the observed open clusters with old ages and large masses, which is also supported by N-body simulations of individual embedded clusters. To explain the mass and radius distributions of the observed open clusters, initial embedded clusters with masses larger than 3000 M⊙ are necessary. However, the upper limit of the embedded cluster sample is less than 1000 M⊙. And only few ATLASGAL clumps have a mass larger than 3000 M⊙. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters. If the Milky Way has a burst-like and time-dependent star formation history, the currently observed open clusters with old ages and large masses may come from massive clumps in the past. There is also a great possibility that these open clusters originate from multiple embedded clusters mergers. We compared the separation of open clusters and the typical size of molecular clouds, and found that most molecular clouds may only form one open cluster, which supports the merger scenario. Further study is necessary to distinguish between different scenarios.
@@
Massive star-forming regions (MSFRs) constitute a primary mode of star formation within the Galaxy (Lada & Lada, 2003; Motte et al., 2018). Young stellar clusters within these regions can potentially evolve into open clusters. It has been proposed that the majority of observed stars, if not all, originate from embedded clusters (Kroupa, 1995a, c; Lada & Lada, 2003; Kroupa, 2005; Megeath et al., 2016; Dinnbier et al., 2022). Despite their significance, the characteristics of embedded clusters and their connection to open clusters remain unclear. The physical processes responsible for the fragmentation of cluster-forming gas and the subsequent dynamics of the resulting stars are not well understood (Megeath et al., 2016). While it is probable that embedded clusters are precursors to open clusters, Lada & Lada (2003) found that only 7% of embedded clusters survive the gas dispersal phase.
@@
Understanding the structure and dynamical evolution of embedded clusters represents a crucial step toward unraveling the formation processes of open clusters. Recent observational advancements indicate that early expansion plays a pivotal role in shaping the development of young star clusters. Utilizing Gaia DR2 and DR3 data alongside multi-object spectroscopy, studies have unveiled the expansion phenomena in very young clusters within star-forming regions through both case studies (Wright et al., 2019; Cantat-Gaudin et al., 2019; Kuhn et al., 2020; Lim et al., 2020; Swiggum et al., 2021; Lim et al., 2022; Mužić et al., 2022; Das et al., 2023; Flaccomio et al., 2023) and statistical works (Kuhn et al., 2019; Della Croce et al., 2023b; Wright et al., 2023). N-body simulations of massive star clusters with monolithic structures and residual gas expulsion (Kroupa et al., 2001; Banerjee & Kroupa, 2013, 2014, 2015; Haghi et al., 2020) as well as semi-analytical models (Dib et al., 2007, 2010; Dib, 2023) have consistently replicated the characteristics of well-observed very young massive clusters such as the Orion Nebula Cluster (ONC), R136, NGC 3603, and the Arches cluster offering insights into their initial formation conditions. For lower-mass embedded clusters, Zhou et al. (2024) compiled various cluster samples to examine the early expansion dynamics using numerical simulation recipes from Kroupa et al. (2001); Banerjee & Kroupa (2013, 2014, 2015); Dinnbier et al. (2022). A special focus of that work was to study the evolution of the initial mass-radius relation of embedded clusters.
@@
In Zhou et al. (2024), direct N-body simulations were employed to model the early expansion of embedded clusters after the expulsion of their residual gas. The models effectively fit the observational data on cluster radii from various sources, indicating that observed very young clusters are likely in an expanding phase. They also suggest that even embedded clusters within ATLASGAL clumps with HII regions may already be undergoing expansion. In Zhou et al. (2024b), the dendrogram algorithm was employed to analyze surface density maps of stars derived from the Massive Young Star-Forming Complex Study in Infrared and X-ray (MYStIX) project (Feigelson et al., 2013; Kuhn et al., 2015b). This approach decomposes the maps into hierarchical structures, revealing the multi-scale features of star clusters. Additionally, the minimum spanning tree (MST) method was utilized to quantify the distances between clusters and clumps within each massive star-forming region. Notably, the distances between clusters, between clumps, and between clusters and clumps were found to be comparable. This suggests that the evolution from clumps to embedded clusters occurs locally and independently, with minimal influence on surrounding objects. In this work, we will further investigate the evolution from embedded clusters to open clusters by analyzing observational data and conducting N-body simulations.
@@ 2.1 Embedded clusters
In Zhou et al. (2024b), we identified a sample of embedded clusters based on the surface density maps of stars for 17 MSFRs with distances less than 3.6 kpc from the MYStIX project. Apart from this sample, we also add the embedded cluster catalogs of Lada et al. (1991); Carpenter et al. (1993); Lada & Lada (2003); Carpenter (2000); Kumar et al. (2006); Faustini et al. (2009), and require the distance of the clusters to be less than 3.6 kpc.
@@ 2.3 Open clusters
Hunt & Reffert (2023) conducted a blind, all-sky search for open clusters using 729 million sources from Gaia DR3 down to magnitude G 20, creating a homogeneous catalog of 7167 clusters. In Table.3 of Hunt & Reffert (2024a), the catalog includes 5647 open clusters (OCs) and 1309 moving groups (MGs), where 3530 OCs and 539 MGs are of high-quality. For the 3530 high-quality open clusters, most of them (3103) have the distance less than 3.6 kpc.
@@ 3.1 Mass and radius
For the ATLASGAL clumps in Urquhart et al. (2022), the radius was determined from the number of pixels within the Full Width at Half Maximum (FWHM) contour, i.e. above 50 per cent of the peak of the ATLASGAL dust continuum emission. The clump’s mass, also measured at FWHM, is determined by integrating the flux density at 870 m within the 50 percent contour. Furthermore, the radii of ATLASGAL clumps with HII regions (HII-clumps) were confirmed to be equal to the half-mass radii of embedded clusters in the clumps in Zhou et al. (2024). In Hunt & Reffert (2024a), the radius of open clusters was determined by including half of the members (r50), close to the half-mass radius. For embedded clusters, their radii have been unified to the half-mass radii, as done in Zhou et al. (2024). Therefore, the radii of clumps, embedded clusters and open clusters can be compared directly. Generally, the radii of open clusters are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters, as presented in Zhou et al. (2024), also clearly shown in Sec.A.
In Fig.1(b), the mass distribution decreases from open clusters to clumps to embedded clusters in sequence. Interestingly, the mass distribution of clumps, when multiplied by 0.33, is comparable to that of embedded clusters, implying that 0.33 could be a typical star formation efficiency (SFE) for clumps in the Milky Way. The mass distribution of the clumps is significantly smaller than that of the open clusters. The mass of ATLASGAL clumps in Urquhart et al. (2018) is the full mass rather than the FWHM mass in Urquhart et al. (2022), which is still significantly smaller than the mass of open clusters.
These offsets raise questions about the origin of open clusters with masses significantly larger than those of their progenitors (clumps or embedded clusters). From clumps to embedded clusters, the SFE is 1, and there is considerable mass loss during the evolution. Therefore, for a single clump to evolve into a single embedded cluster and then evolve into a single open cluster, with the current clump mass distribution in the Milky Way, this evolutionary sequence cannot produce the observed open clusters’ mass distribution, as verified by the direct N-body simulations in Sec.3.3.
If an open cluster originates from an embedded cluster, the initial mass of the embedded cluster should be much larger than the open cluster, due to the considerable mass loss during the evolution, as shown in Sec.3.4. However, the opposite is seen in Fig.1(b), where the mass distribution of open clusters is significantly larger than that of embedded clusters (about one order of magnitude), suggesting that most open clusters do not originate from single embedded clusters.
@@ 3.2 Separations of open clusters
We used the minimum spanning tree (MST) analysis to measure the inter-cluster spacing based on the 3D coordinates of open clusters. The MST will connect all clusters in a network with the least amount of total distance, without creating any loops. The length of the line connecting two clusters represents the separation between them. In Fig.2, the typical inter-cluster separation of 3103 high-quality open clusters is 80 pc. Considering all 5858 open clusters and moving groups (¡ 3.6 kpc), the lower limit of the separation is 60 pc. However,the most probable size for molecular clouds in the entire Milky Way disk is 30 pc (Miville-Deschênes et al., 2017). This implies that most molecular clouds may only form one open cluster. Each molecular cloud typically includes many embedded clusters, as presented in Sec.A. If each individual embedded cluster could evolve into an open cluster, the spacing between open clusters should not be so large, at least it should be smaller than the size of the parental molecular cloud.
In Zhou et al. (2024b), for the MSFRs in the MYStIX project, the separations between clusters, between clumps, and between clusters and clumps are quite comparable. The typical separation is 1 pc. Therefore, as the embedded clusters in each MSFR expand, they are able to undergo mergers. Significant differences in separation for embedded clusters and open clusters indicate that there should be considerable mergers between embedded clusters during the process of evolving into open clusters. We will discuss this picture in a forthcoming work.
@@ 3.3.1 Fast gas expulsion
The parameters for the simulation are summarized from previous works, i.e. Kroupa et al. (2001); Baumgardt & Kroupa (2007); Banerjee et al. (2012); Banerjee & Kroupa (2013, 2014, 2015); Oh et al. (2015); Oh & Kroupa (2016); Banerjee & Kroupa (2017); Brinkmann et al. (2017); Oh & Kroupa (2018); Wang et al. (2019); Pavlik et al. (2019); Dinnbier et al. (2022); Zhou et al. (2024). The previous simulations have already demonstrated the effectiveness and rationality of the parameter settings (see below). The influence of different parameter settings on simulation results and the discussion of multi-dimensional parameter space can also be found in the aforementioned literature. In this work, we mainly utilize the mature simulation recipe to interpret observational data. The model parameters of the simulations are listed in Table.LABEL:para.
@@
We computed three clusters with the masses in stars of 300 M⊙, 1000 M⊙ and 3000 M⊙ for the first 100 Myr. The initial density profile of all clusters is the Plummer profile (Aarseth et al., 1974; Heggie & Hut, 2003; Kroupa, 2008). A Plummer would be an appropriate initial profile since molecular clouds’ filaments in which stars form have been found to have Plummer-like cross sections (Malinen et al., 2012; André et al., 2022), and open star clusters can also be described by the Plummer model (Röser et al., 2011; Röser & Schilbach, 2019). Moreover, such a specific initial profile does not significantly affect the overall expansion rate of a cluster, as discussed in Banerjee & Kroupa (2017), which is primarily governed by the total stellar mass loss and the dynamical interactions occurring within the inner part of the cluster. The half-mass radius of the cluster is given by the relation (Marks & Kroupa, 2012), i.e.
@@
All clusters are fully mass segregated (=1), no fractalization, and in the state of virial equilibrium (=0.5). and are the degree of mass segregation and the virial ratio of the cluster, respectively. More details can be found in Küpper et al. (2011) and the user manual for the McLuster code. The initial segregated state is detected for young clusters and star-froming clumps/clouds (Littlefair et al., 2003; Chen et al., 2007; Portegies Zwart et al., 2010; Kirk & Myers, 2011; Getman et al., 2014; Lane et al., 2016; Alfaro & Román-Zúñiga, 2018; Plunkett et al., 2018; Pavlik et al., 2019; Nony et al., 2021; Zhang et al., 2022; Xu et al., 2024), but the degree of mass segregation is not clear. In simulations of the very young massive clusters R136 and NGC 3603 with gas expulsion by Banerjee & Kroupa (2013), mass segregation does not influence the results. In Zhou et al. (2024), we compared =1 (fully mass segregated) and =0.5 (partly mass segregated), and found the results are similar. We also discussed settings with and without fractalization in Zhou et al. (2024), the results of both are also consistent. The IMFs of the clusters are chosen to be canonical (Kroupa, 2001) with the most massive star following the relation of Weidner et al. (2013), i.e.
@@
where = , = , = -0.66, = 1.08, = -0.150, and = 0.0084. We assume the clusters to be at solar metallicity, i.e. (von Steiger & Zurbuchen, 2016). The clusters traverse circular orbits within the Galaxy, positioned at a Galactocentric distance of 8.5 kpc, moving at a speed of 220 km s-1.
@@
The initial binary setup follows the method described in Wang et al. (2019). All stars are initially in binaries, i.e. =1, is the primordial binary fraction. Kroupa (1995b, d) propose that stars with masses below a few are initially formed with a universal binary distribution function and that star clusters start with a binary fraction. Inverse dynamical population synthesis was employed to derive the initial distributions of binary periods and mass ratios. Belloni et al. (2017) introduces an updated model of pre-main-sequence eigenevolution, originally developed by Kroupa (1995d), to account for the observed correlations between mass ratio, period, and eccentricity in short-period systems. For low-mass binaries, we adopt the formalism developed in Kroupa (1995b, d); Belloni et al. (2017) to characterize the period, mass ratio and eccentricity distributions. For high-mass binaries (OB stars with a mass M⊙), we utilize the Sana et al. (2012) distribution, which is derived from O stars in open clusters. The distribution functions of the period, mass ratio and eccentricity are presented in Oh et al. (2015); Belloni et al. (2017).
@@
Accurately modeling gas removal from embedded clusters is challenging due to the complexity of radiation hydrodynamical processes, which involve uncertain and intricate physical mechanisms. To simplify the approach, we simulate the key dynamical effects of gas expulsion by applying a diluting, spherically symmetric external gravitational potential to a model cluster, following the method presented in Kroupa et al. (2001); Banerjee & Kroupa (2013). This analytical approach is partially validated by Geyer & Burkert (2001), who conducted comparison simulations using the SPH (smoothed particle hydrodynamics) method to treat the gas. Specifically, we use the potential of the spherically-symmetric, time-varying mass distribution
@@
Here, is the total mass in gas which is spatially distributed with the Plummer density distribution (Kroupa, 2008) and starts depleting with timescale after a delay of . The Plummer radius of the gas distribution is kept time-invariant (Kroupa et al., 2001). This assumption is an approximate model of the effective gas reduction within the cluster in the situation that gas is blown out while new gas is also accreting into the cluster along filaments such that the gas mass ends up being reduced with time but the radius of the gas distribution remains roughly constant. As discussed in Urquhart et al. (2022), the mass and radius distributions of the ATLASGAL clumps at different evolutionary stages are quite comparable. We use an average velocity of the outflowing gas of km s-1, which is the typical sound-speed in an HII region. This gives , where is the initial half-mass radius of the cluster. As for the delay-time, we take the representative value of Myr (Kroupa et al., 2001), this being about the life-time of the ultra-compact HII region. As shown in Banerjee & Kroupa (2013), varying the delay time, , primarily results in a temporal shift in the cluster’s rapid expansion phase, without significantly impacting its subsequent evolution for times greater than . Protoclusters typically form in hub-filament systems (Motte et al., 2018; Vázquez-Semadeni et al., 2019; Kumar et al., 2020; Zhou et al., 2022), which are located in hub regions. Compared to the surrounding filamentary gas structures, the hub region, as the center of gravitational collapse, is usually more regular, as shown in Zhou et al. (2022, 2024). Thus, modeling a spherically-symmetric mass distribution is appropriate, see Sec.3.3.3 for more discussion.
@@
In this work, we assume a star formation efficiency SFE 0.33 as a benchmark, i.e. = 2. This value has been widely used in previous simulations listed above and has been proven to be effective in reproducing the observational properties of stellar clusters. Such a SFE is also consistent with the value obtained from hydrodynamical calculations including self-regulation (Machida & Matsumoto, 2012; Bate et al., 2014) and as well with observations of embedded systems in the solar neighborhood (Lada & Lada, 2003; Megeath et al., 2016). In Zhou et al. (2024b), by comparing the mass functions of the ATLASGAL clumps and the identified embedded clusters, we found that a star formation efficiency of 0.33 can be a typical value for the ATLASGAL clumps, also presented in Fig.1(b). In Zhou et al. (2024a), assuming SFE = 0.33, it was shown that the total bolometric luminosity of synthetic embedded clusters can well fit the observed bolometric luminosity of ATLASGAL clumps with HII regions. In Zhou et al. (2024), we directly calculated the SFE of ATLASGAL clumps with HII regions, and found a median value of 0.3.

### v2

@@
22email: sami.dib@gmail.com 33institutetext: Helmholtz-Institut für Strahlen- und Kernphysik (HISKP), Universität Bonn, Nussallee 14–16, 53115 Bonn, Germany
33email: pkroupa@uni-bonn.de 44institutetext: Charles University in Prague, Faculty of Mathematics and Physics, Astronomical Institute, V Holešovičkách 2, CZ-180 00 Praha 8, Czech Republic
@@
We categorized clumps, embedded clusters, and open clusters and conducted a comparative analysis of their physical properties. Overall, the radii of open clusters are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters. The open clusters have significantly higher masses than embedded clusters, by about one order of magnitude. Given the current mass distribution of clumps in the Milky Way, the evolutionary sequence from a single clump evolving into an embedded cluster and subsequently into an open cluster cannot account for the observed open clusters with old ages and high masses, which is also supported by N-body simulations of individual embedded clusters. To explain the mass and radius distributions of the observed open clusters, initial embedded clusters with masses higher than 3000 M⊙ are necessary. However, the upper limit of the embedded cluster sample is less than 1000 M⊙, and only a few ATLASGAL clumps have a mass higher than 3000 M⊙. Thus, the currently observed clumps cannot be the ”direct” precursors of the currently observed open clusters. If the Milky Way has a burst-like and time-dependent star formation history, the currently observed open clusters with old ages and high masses may come from massive clumps in the past. There is also a very real possibility that these open clusters originate from post-gas expulsion coalescence of multiple embedded clusters. We compared the separation of open clusters and the typical size of molecular clouds, and find that most molecular clouds may only form one open cluster, which supports the scenario of post-gas expulsion coalescence. Further study is necessary to distinguish between the different scenarios.
@@
Massive star-forming regions (MSFRs) constitute a primary mode of star formation within the Galaxy (Lada & Lada, 2003; Motte et al., 2018). Young stellar clusters within these regions can potentially evolve into open clusters (OCs). It has been proposed that the majority of observed stars, if not all, originate from embedded clusters (Kroupa, 1995a, c; Lada & Lada, 2003; Kroupa, 2005; Megeath et al., 2016; Dinnbier et al., 2022). Despite their significance, the characteristics of embedded clusters and their connection to OCs remain unclear. The physical processes responsible for the fragmentation of cluster-forming gas and the subsequent dynamics of the resulting stars are not well understood (Megeath et al., 2016). While it is probable that embedded clusters are precursors to OCs, Lada & Lada (2003) found that only 7% of embedded clusters survive the gas dispersal phase.
@@
Understanding the structure and dynamical evolution of embedded clusters represents a crucial step toward unraveling the formation processes of OCs. Recent observational advancements indicate that early expansion plays a pivotal role in shaping the development of young star clusters. Utilizing Gaia Data Release (DR) 2 and DR3 data alongside multi-object spectroscopy, studies have unveiled the expansion phenomena in very young clusters within star-forming regions through both case studies (Wright et al., 2019; Cantat-Gaudin et al., 2019; Kuhn et al., 2020; Lim et al., 2020; Swiggum et al., 2021; Lim et al., 2022; Mužić et al., 2022; Das et al., 2023; Flaccomio et al., 2023) and statistical works (Kuhn et al., 2019; Della Croce et al., 2023b; Wright et al., 2023). N-body simulations of massive star clusters with monolithic structures and residual gas expulsion (Kroupa et al., 2001; Banerjee & Kroupa, 2013, 2014, 2015; Haghi et al., 2020) as well as semi-analytical models (Dib et al., 2007, 2010; Dib, 2023) have consistently replicated the characteristics of well-observed very young massive clusters such as the Orion Nebula Cluster, R136, NGC 3603, and the Arches cluster, offering insights into their initial formation conditions. For lower-mass embedded clusters, Zhou et al. (2024) compiled various cluster samples to examine the early expansion dynamics using numerical simulation recipes from Kroupa et al. (2001), Banerjee & Kroupa (2013, 2014, 2015), and Dinnbier et al. (2022). A special focus of that work was the evolution of the initial mass-radius relation of embedded clusters.
@@
In Zhou et al. (2024), direct N-body simulations were employed to model the early expansion of embedded clusters after the expulsion of their residual gas. The models effectively fit the observational data on cluster radii from various sources, indicating that observed very young clusters are likely in an expanding phase. They also suggest that even embedded clusters within ATLASGAL clumps with HII regions may already be undergoing expansion. In Zhou et al. (2024b), the dendrogram algorithm was employed to analyze surface density maps of stars derived from the Massive Young Star-Forming Complex Study in Infrared and X-ray (MYStIX) project (Feigelson et al., 2013; Kuhn et al., 2015b). This approach decomposes the maps into hierarchical structures, revealing the multi-scale features of star clusters. Additionally, the minimum spanning tree (MST) method was utilized to quantify the distances between clusters and clumps within each MSFR. Notably, the distances between clusters, between clumps, and between clusters and clumps were found to be comparable. This suggests that the evolution from clumps to embedded clusters occurs locally and independently, with minimal influence on surrounding objects. In this work we further investigated the evolution from embedded clusters to OCs by analyzing observational data and conducting N-body simulations.
@@ 2.1 Embedded clusters
In Zhou et al. (2024b) we identified a sample of embedded clusters based on the surface density maps of stars for 17 MSFRs with distances less than 3.6 kpc from the MYStIX project. To this we added the embedded cluster catalogs of Lada et al. (1991), Carpenter et al. (1993), Lada & Lada (2003), Carpenter (2000), Kumar et al. (2006), and Faustini et al. (2009) and required the distance of the clusters to be less than 3.6 kpc.
@@ 2.3 Open clusters
Hunt & Reffert (2023) conducted a blind, all-sky search for OCs using 729 million sources from Gaia DR3 down to magnitude G 20, creating a homogeneous catalog of 7167 clusters. The Hunt & Reffert (2024a) catalog (see their Table 3) includes 5647 OCs and 1309 moving groups, of which 3530 OCs and 539 moving groups are of high quality. For the 3530 high-quality OCs, most of them (3103) have a distance of less than 3.6 kpc.
@@ 3.1 Mass and radius
For the ATLASGAL clumps in Urquhart et al. (2022), the radius was determined from the number of pixels within the full width at half maximum (FWHM) contour (i.e., above 50 percent of the peak of the ATLASGAL dust continuum emission). The clump’s mass, also measured at the FWHM, was determined by integrating the flux density at 870 m within the 50 percent contour. Furthermore, the radii of ATLASGAL clumps with HII regions (HII clumps) were confirmed to be equal to the half-mass radii of embedded clusters in the clumps in Zhou et al. (2024). In Hunt & Reffert (2024a), the radius of OCs was determined by including half of the members (r50) that is close to the half-mass radius. For embedded clusters, their radii were unified to the half-mass radii, as done in Zhou et al. (2024). Therefore, the radii of clumps, embedded clusters, and OCs can be directly compared. Generally, the radii of OCs are significantly larger than those of embedded clusters and clumps. The radii of embedded clusters are larger than those of clumps, which may be due to the expansion of embedded clusters, as presented in Zhou et al. (2024) and clearly shown in Appendix A.
In Fig.1(b), the mass distribution decreases from OCs to clumps to embedded clusters. Interestingly, the mass distribution of clumps, when multiplied by 0.33, is comparable to that of embedded clusters, implying that 0.33 could be a typical star formation efficiency (SFE) for clumps in the Milky Way. The mass distribution of the clumps is significantly smaller than that of the OCs. The mass of ATLASGAL clumps in Urquhart et al. (2018) is the full mass rather than the FWHM mass in Urquhart et al. (2022), which is still significantly lower than the mass of OCs.
These offsets raise questions about the origin of OCs with masses significantly higher than those of their progenitors (clumps or embedded clusters). From clumps to embedded clusters, the SFE is 0.3 (Lada & Lada, 2003; Megeath et al., 2016; Zhou et al., 2024b, a, 2024), and there is considerable mass loss during the evolution. Therefore, for a single clump to evolve into a single embedded cluster and then evolve into a single OC, with the current clump mass distribution in the Milky Way, this evolutionary sequence cannot produce the observed OCs’ mass distribution, as verified by the direct N-body simulations in Sect. 3.3.
If an OC originates from an embedded cluster, the initial mass of the embedded cluster should be much higher than that of the OC due to the considerable mass loss during the evolution, as shown in Sect. 3.4. However, the opposite is seen in Fig.1(b), where the mass distribution of OCs is significantly larger than that of embedded clusters (by about one order of magnitude), suggesting that most OCs do not originate from single embedded clusters.
@@ 3.2 Separations of open clusters
We used the minimum spanning tree (MST) analysis to measure the inter-cluster spacing based on the 3D coordinates of OCs. The MST connects all clusters in a network with the least amount of total distance, without creating any loops. The length of the line connecting two clusters represents the separation between them. In Fig. 2 the typical inter-cluster separation of 3103 high-quality OCs is 80 pc. Considering all 5858 OCs and moving groups (¡ 3.6 kpc), the lower limit of the separation is 60 pc. However, the most probable size for molecular clouds in the entire Milky Way disk is 30 pc (Miville-Deschênes et al., 2017). This implies that most molecular clouds may only form one OC. Each molecular cloud typically includes many embedded clusters, as presented in Appendix A. If each individual embedded cluster could evolve into an OC, the spacing between OCs should not be so large; it should at least be smaller than the size of the parental molecular cloud.
In Zhou et al. (2024b), for the MSFRs in the MYStIX project, the separations between clusters, between clumps, and between clusters and clumps are quite comparable. The typical separation is 1 pc. Therefore, as the embedded clusters in each MSFR expand, they are able to undergo merger (post-gas expulsion coalescence). Significant differences in separation for embedded clusters and OCs indicate that there are considerable mergers between embedded clusters as they evolve into OCs. We will discuss this picture in a forthcoming work.
@@ 3.3 Single embedded cluster simulation
Notes: and are the mass and the half-mass radius of the cluster. is the mass of the most massive star in the cluster. and are the decay time and the delay time during the gas expulsion process. The tree time step (-s), the changeover radius (-r), the neighbor searching radius (–r-search-min), and the multiple group radius (–r-bin) are the important parameters in PeTar.
@@ 3.3.1 Fast gas expulsion
The parameters for the simulation are summarized from previous works (i.e., Kroupa et al. 2001; Baumgardt & Kroupa 2007; Banerjee et al. 2012; Banerjee & Kroupa 2013, 2014, 2015; Oh et al. 2015; Oh & Kroupa 2016; Banerjee & Kroupa 2017; Brinkmann et al. 2017; Oh & Kroupa 2018; Wang et al. 2019; Pavlik et al. 2019; Dinnbier et al. 2022; Zhou et al. 2024). The previous simulations have already demonstrated the effectiveness and rationality of the parameter settings (see below). The influence of different parameter settings on simulation results and the discussion of the multidimensional parameter space can also be found in the works cited above. In this work, we mainly utilized the mature simulation recipe to interpret observational data. The model parameters of the simulations are listed in Table.LABEL:para.
@@
We computed three clusters with star masses of 300 M⊙, 1000 M⊙, and 3000 M⊙ in the first 100 Myr. The initial density profile of all clusters was the Plummer profile (Aarseth et al., 1974; Heggie & Hut, 2003; Kroupa, 2008), an appropriate choice since the molecular clouds’ filaments in which stars form have been found to have Plummer-like cross sections (Malinen et al., 2012; André et al., 2022), and open star clusters can also be described by the Plummer model (Röser et al., 2011; Röser & Schilbach, 2019). Moreover, such a specific initial profile does not significantly affect the overall expansion rate of a cluster, as discussed in Banerjee & Kroupa (2017), which is primarily governed by the total stellar mass loss and the dynamical interactions occurring within the inner part of the cluster. The half-mass radius, , of the cluster is given by the relation (Marks & Kroupa, 2012):
@@
All clusters are fully mass segregated (=1), with no fractalization, and in a state of virial equilibrium (=0.5). and are the degree of mass segregation and the virial ratio of the cluster, respectively. More details can be found in Küpper et al. (2011) and the user manual for the McLuster code. The initial segregated state is detected for young clusters and star-forming clumps/clouds (Littlefair et al., 2003; Chen et al., 2007; Portegies Zwart et al., 2010; Kirk & Myers, 2011; Getman et al., 2014; Lane et al., 2016; Alfaro & Román-Zúñiga, 2018; Plunkett et al., 2018; Pavlik et al., 2019; Nony et al., 2021; Zhang et al., 2022; Xu et al., 2024), but the degree of mass segregation is not clear. In simulations of the very young massive clusters R136 and NGC 3603 with gas expulsion by Banerjee & Kroupa (2013), mass segregation does not influence the results. In Zhou et al. (2024), we compared =1 (fully mass segregated) and =0.5 (partly mass segregated) and found similar results. We also discussed settings with and without fractalization in Zhou et al. (2024); the results of the two are also consistent. The initial mass functions of the clusters were chosen to be canonical (Kroupa, 2001), with the most massive star following the relation of Weidner et al. (2013):
@@
where = , = , = -0.66, = 1.08, = -0.150, and = 0.0084. We assumed the clusters to be at solar metallicity (i.e., ; von Steiger & Zurbuchen, 2016). The clusters travel along circular orbits within the Galaxy, positioned at a galactocentric distance of 8.5 kpc, and are moving at a speed of 220 km s-1.
@@
The initial binary setup follows the method described in Wang et al. (2019). All stars are initially in binaries, that is to say, =1, where is the primordial binary fraction. Kroupa (1995b, d) propose that stars with masses below a few are initially formed with a universal binary distribution function and that star clusters start with a binary fraction. Inverse dynamical population synthesis was employed to derive the initial distributions of binary periods and mass ratios. Belloni et al. (2017) introduce an updated model of pre-main-sequence eigenevolution, originally developed by Kroupa (1995d), to account for the observed correlations between the mass ratio, period, and eccentricity in short-period systems. For low-mass binaries, we adopted the formalism developed in Kroupa (1995b, d) and Belloni et al. (2017) to characterize the period, mass ratio, and eccentricity distributions. For high-mass binaries (OB stars with masses M⊙), we utilized the Sana et al. (2012) distribution, which is derived from O stars in OCs. The distribution functions of the period, mass ratio, and eccentricity are presented in Oh et al. (2015) and Belloni et al. (2017).
@@
Accurately modeling gas removal from embedded clusters is challenging due to the complexity of radiation hydrodynamical processes, which involve uncertain and intricate physical mechanisms. To simplify the approach, we simulated the key dynamical effects of gas expulsion by applying a diluting, spherically symmetric external gravitational potential to a model cluster, following the method presented in Kroupa et al. (2001) and Banerjee & Kroupa (2013). This analytical approach is partially validated by Geyer & Burkert (2001), who conducted comparison simulations using the smoothed particle hydrodynamics method to treat the gas. The hydrodynamics+N-body simulations in Farias2024-527 also confirm that the exponential decay model presented in Eq. 3 generally provides a good description of gas removal driven by radiation and wind feedback. Specifically, we used the potential of the spherically symmetric, time-varying mass distribution:
@@
Here, is the total mass of the gas; it is spatially distributed with the Plummer density distribution (Kroupa, 2008) and starts depleting after a delay of , and is totally depleted within a timescale of . The Plummer radius of the gas distribution is kept time-invariant (Kroupa et al., 2001). This assumption is an approximate model of the effective gas reduction within the cluster in the situation that gas is blown out while new gas is also accreting into the cluster along filaments such that the gas mass ends up being reduced with time but the radius of the gas distribution remains roughly constant. As discussed in Urquhart et al. (2022), the mass and radius distributions of the ATLASGAL clumps at different evolutionary stages are quite comparable. We used an average velocity of the outflowing gas of km s-1, which is the typical sound speed in an HII region. This gives , where is the initial half-mass radius of the cluster. As for the delay time, we took the representative value of Myr (Kroupa et al., 2001), this being about the lifetime of the ultracompact HII region. As shown in Banerjee & Kroupa (2013), varying the delay time, , primarily results in a temporal shift in the cluster’s rapid expansion phase, without significantly impacting its subsequent evolution for times greater than . Protoclusters typically form in hub-filament systems (Motte et al., 2018; Vázquez-Semadeni et al., 2019; Kumar et al., 2020; Zhou et al., 2022), which are located in hub regions. Compared to the surrounding filamentary gas structures, the hub region, as the center of gravitational collapse, is usually more regular, as shown in Zhou et al. (2022, 2024). Thus, modeling a spherically symmetric mass distribution is appropriate (see Sect. 3.3.3 for more discussion).
@@
In this work, we assumed a SFE 0.33 as a benchmark (i.e., = 2. This value has been widely used in the simulations cited above and has been proven effective in reproducing the observational properties of stellar clusters. Such a SFE is also consistent with the value obtained from hydrodynamical calculations that include self-regulation (Machida & Matsumoto, 2012; Bate et al., 2014) and as well with observations of embedded systems in the solar neighborhood (Lada & Lada, 2003; Megeath et al., 2016). In Zhou et al. (2024b), by comparing the mass functions of the ATLASGAL clumps and the identified embedded clusters, we found that a SFE of 0.33 is typical for the ATLASGAL clumps, which are also presented in Fig.1(b) here. In Zhou et al. (2024a), assuming SFE = 0.33, it was shown that the total bolometric luminosity of synthetic embedded clusters can fit the observed bolometric luminosity of ATLASGAL clumps with HII regions. In Zhou et al. (2024), we directly calculated the SFE of ATLASGAL clumps with HII regions and found a median value of 0.3.